# GOFA: A Generative One-For-All Model for Joint Graph Language Modeling

### v1

@@ F.3 Zero-shot learning
Model weight_decay gradient_accum llm_max_length train_sample_size
GOFA-Llama2-Node 0.1 4 256 10000
GOFA-Mistral-Node 0.1 64 256 10000/40000
GOFA-Llama2-Link 0.0001 4 256 60000
GOFA-Mistral-Link 0.1 64 256 40000
Detail of GOFA. For the GOFA, we fine-tune the model from the pre-training checkpoint. In fine-tuning, we will train the parameters of GNN and LoRA layers in the LLM decoder. To comprehensively evaluate the performance of GOFA, We separately fine-tune the GOFA on different datasets. Specifically, we design two different settings. In the first setting, we fine-tune the model using the Arxiv dataset with the node classification task. In the second setting, we fine-tune the model using the PubMed dataset with the link prediction task. For each setting, we fine-tune both the Llama2-7B and Mistral-7B versions, which results in 4 different versions. We denote as GOFA-Llama2-Node, GOFA-Mistral-Node, GOFA-Llama2-Link, GOFA-Mistral-Link, and GOFA-Mistral-Joint, respectively. For all evaluation and pre-training datasets, we design multiple prompt templates with instructions to let the model select the correct label from the provided label list. For each label in each dataset, we use the GPT-4 to generate a short description for the label. The detailed prompt examples for all datasets are shown in Table 9. For fine-tuning on Arixv, since it is hard to include all 40 ways in the prompt, we randomly sampled 5 ways during the training for each sample. For each pre-training dataset, we randomly sample a fixed number of training samples in a stratified way. The detailed parameters for fine-tuning are listed in Table 10. All parameters not listed in the table are the same as the pre-training setting. For all training versions, we directly evaluate the model on the test set of all evaluation datasets. We evaluate the model on the whole test set except Products and FB15K237, which only sample 10000 samples for evaluation. For evaluation, we will match the text output generated by the GOFA with the ground true label to compute the accuracy of the classification task. For the regression task, we will extract the number from the output text and compute the metric with the correct value.
F.4 Supervised-learning
@@
0.0001 0.1 0.5 4 256
Setting. For the supervised-learning setting, we select Cora (node/link), PubMed (node/link), Arxiv, WikiCS, WN18RR, FB15K237, Products, and ML1M datasets for the evaluation. For all datasets, we utilize the default split described in Appendix C. To ensure a fair comparison, we employ subgraph sampling for GOFA and all baseline methods. For all datasets, the sampling hop is 3 and the maximum nodes per hop are 5, except for the ML1M, which is 1 and 15 respectively.
Detail of baselines. For the traditional GNN methods, we include GCN Kipf and Welling [2017] and GAT Veličković et al. [2018]. To ensure a fair comparison, we use Llama2-7B to convert raw texts in all datasets to sentence embedding and use this as the model’s input node/edge features. We re-implement both methods in order to adapt the original method with subgraph input. Specifically, for but node/link-level tasks, we will add labeling trick Zhang et al. [2021] to the target nodes at the beginning. After message passing, we will use the summation pooling on all target nodes and use the result embedding for the prediction. For traditional GNN methods, we train and evaluate each dataset independently. For all datasets, we search the number of layers and dropout parameters. For each parameter set, we repeat the experiment 4 times select the parameter set with the best validation performance, and report the performance on the test set. For the graph foundation model, we include OFA Liu et al. [2023a] as the baseline. The OFA is simultaneously trained and evaluated on all datasets. To ensure a fair comparison, we get their code from the original source and train the model on Cora (node/link), PubMed (node/link), Arxiv, WikiCS, WN18RR, and FB15K237 dataset using the Llama2-7b as base LLM model. Similarly, for OFA, we use the same subgraph sampling parameters as all other methods. For other parameters, we use the default parameter provided in their code. We only run the model one time and report the final performance.
@@
ML1M This is a recommendation graph from a movie rating platform. Nodes represent user or movie in the platform and edges represent the rating a user gives to a movie. Please predict the user taste to this movie, ranging from 1 to 5.
Detail of GOFA. For the GOFA, we fine-tune the model from the pre-training checkpoint on both Llama2 and Mistral. In fine-tuning, we will train the parameters of GNN and LoRA layers in the LLM decoder. We simultaneously fine-tune the model on the train set of Cora-node, Cora-link, PubMed-node, PubMed-link, Arxiv, WikiCS, WN18RR, FB15K237, Products, and ML1M. For each dataset, we will randomly sample a fixed number of training samples for each epoch with stratified sampling. The sample number for each dataset is 1200, 15000, 1300, 20000, 16000, 8000, 20000, 20000, 10000, 15000, respectively. We fine-tune the model for 1 epochs. The detailed parameters for fine-tuning are listed in Table 11. For each dataset, we create a prompt for the LLM decoder to generate the desired answer. In a supervised setting, we ask the LLM model directly to generate the correct answer, instead of doing the selection from the given list. The detailed prompt for each dataset is listed in Table 12. For evaluation, we will match the text output generated by the GOFA with the ground true label to compute the accuracy of the classification task. For the regression task, we will extract the number from the output text and compute the metric with the correct value.

### v2

@@ F.4 Zero-shot learning
lr weight_decay gradient_accum llm_max_length
0.0001 0.1 64 256
Detail of GOFA. For the GOFA, we fine-tune the model from the pre-training checkpoint. In fine-tuning, we will train the parameters of GNN and LoRA layers in the LLM decoder. To comprehensively evaluate the performance of GOFA, We separately fine-tune the GOFA on different datasets. Specifically, we design two different settings. In the first setting, we fine-tune the model using the Arxiv and Pubmed datasets with both the node classification and link prediction tasks. In the second setting, we add mag240m and Wikikg90m additionally. We denote GOFA-T and GOFA-F, respectively. For GOFA-T, we sample 40000, 80000, 10000, 10000 for Arxiv_link, Arxiv_node, Pubmed_link, and Pubmed_node, respectively. For GOFA-F, we sample 10000, 10000, 40000, 50000, 10000, 10000 for MAG240M, Wikikg90m, Arxiv_link, Arxiv_node, Pubmed_node, and Pubmed_link, respectively. For all evaluation and pre-training datasets, we design multiple prompt templates with instructions to let the model select the correct label from the provided label list. For each label in each dataset, we use the GPT-4 to generate a short description for the label. The detailed prompt examples for all datasets are shown in Table 13 and Table 14. For all MAG240M, Wikikg90m, and Arxiv, since it is hard to include all ways in the prompt, we randomly sampled 10 ways during the training for each sample. For each pre-training dataset, we randomly sample a fixed number of training samples in a random way. The detailed parameters for fine-tuning are listed in Table 15. All parameters not listed in the table are the same as the pre-training setting. For all training versions, we directly evaluate the model on the test set of all evaluation datasets. We evaluate the model on the test set. For datasets with less than 15000 test samples, we evaluate on the whole set. Otherwise, we only randomly select 15000 samples for evaluation, due to the time constraint. For evaluation, we will match the text output generated by the GOFA with the ground true label to compute the accuracy of the classification task. For the regression task, we will extract the number from the output text and compute the metric with the correct value.
F.5 Supervised-learning
@@
0.0001 0.1 0.5 32 256
Setting. For the supervised-learning setting, we select Cora (node/link), PubMed (node/link), Arxiv, WikiCS, WN18RR, FB15K237, and Products datasets for the evaluation. For all datasets, we utilize the default split described in Appendix C. To ensure a fair comparison, we employ subgraph sampling for GOFA and all baseline methods. For all datasets, the sampling hop is 3 and the maximum nodes per hop are 5.
Detail of baselines. For the traditional GNN methods, we include GCN (Kipf & Welling, 2017) and GAT (Veličković et al., 2018). To ensure a fair comparison, we use Llama2-7B to convert raw texts in all datasets to sentence embedding and use this as the model’s input node/edge features. We re-implement both methods in order to adapt the original method with subgraph input. Specifically, for but node/link-level tasks, we will add labeling trick (Zhang et al., 2021) to the target nodes at the beginning. After message passing, we will use the summation pooling on all target nodes and use the result embedding for the prediction. For traditional GNN methods, we train and evaluate each dataset independently. For all datasets, we search the number of layers and dropout parameters. For each parameter set, we repeat the experiment 4 times select the parameter set with the best validation performance, and report the performance on the test set. For constrastive learning methods, we include DGI (Veličković et al., 2018) and BGRL (Thakoor et al., 2021). We directly report results from UniGraph (He & Hooi, 2024) . For the graph foundation model, we include OFA (Liu et al., 2023a) and UniGraph (He & Hooi, 2024) as the baseline. The OFA is simultaneously trained and evaluated on all datasets. To ensure a fair comparison, we get their code from the original source and train the model on Cora (node/link), PubMed (node/link), Arxiv, WikiCS, WN18RR, and FB15K237 dataset using the Llama2-7b as base LLM model. Similarly, for OFA, we use the same subgraph sampling parameters as all other methods. For other parameters, we use the default parameter provided in their code. We only run the model one time and report the final performance. For UniGraph, we directly report results from their original paper.
@@
Detail of GOFA. For the GOFA, we fine-tune the model from the pre-training checkpoint. In fine-tuning, we will train the parameters of GNN and LoRA layers in the LLM decoder. We simultaneously fine-tune the model on the train set of Cora-node, Cora-link, PubMed-node, PubMed-link, Arxiv, WikiCS, WN18RR, FB15K237, and Products. For each dataset, we will randomly sample a fixed number of training samples for each epoch with random sampling. The sample numbers for each dataset is 3000, 40000, 3000, 80000, 105000, 12000, 60000, 120000, and 38000, respectively. We fine-tune the model for 1 epochs. The detailed parameters for fine-tuning are listed in Table 16. For each dataset, we create a prompt for the LLM decoder to generate the desired answer. In a supervised setting, we ask the LLM model directly to generate the correct answer, instead of doing the selection from the given list. The detailed prompt for each dataset is listed in Table 17. For evaluation, we will match the text output generated by the GOFA with the ground true label to compute the accuracy of the classification task. We evaluate the model on the test set. For datasets with less than 15000 test samples, we evaluate on the whole set. Otherwise, we only randomly select 15000 samples for evaluation, due to the time constraint.